March 6, 2023
Return to the calendar.
RC Meeting Information
Current RC: Iuliia Skorodumina.
No RC meeting today.
Sunday Day (03/05). Smooth data taking at 60 uA was interrupted around 3pm by SHMS Q1 interlock (https://logbooks.jlab.org/entry/4143596). Joe Beaufait and Heidi were contacted, they diagnosed the issue as a problem with He level meter and resolved it by the end of the shift (https://logbooks.jlab.org/entry/4143612). Once the magnet was cycled (https://logbooks.jlab.org/entry/4143618), we were ready to go back to production, but by then the beam was gone. After the He level meter fix, the He level dropped down a bit into the yellow alarm region (https://logbooks.jlab.org/entry/4143653), but soon went back to normal. Day Shift summary: https://logbooks.jlab.org/entry/4143548 .
Sunday Swing (03/05). Shift started with no beam due to the issues with a cavity (https://logbooks.jlab.org/entry/4143626), which involved the LCW leak (https://logbooks.jlab.org/entry/4143666). The beam returned around 7pm and we resumed production. Swing Shift summary: https://logbooks.jlab.org/entry/4143633 .
Monday Owl (03/06). Smooth data taking overnight at 60 uA. Owl Shift summary: https://logbooks.jlab.org/entry/4143776 .
For the current setting (Pm=800 MeV/c), we have collected a total of 13620 mC integrated charge out of the planned 15696 mC, which corresponds to 87%.
There was a change of plan and the repair of the 4th pass separator will be performed today during Day shift, followed by the Accelerator recovery during Swing. So, today is the Maintenance Day (https://logbooks.jlab.org/entry/4143879). Let me know if you would like to go down to the Hall.
Work in the Hall
Today is the Maintenance Day. Let me know if you would like to access the Hall.
Current Run Plan
The plan is to continue taking production data for the Pm=800 MeV/c at the highest possible current.